∗ K. Abe,^{(19)} K. Abe,^{(30)} T. Akagi,^{(28)} N.J. Allen,^{(4)} W.W. Ash,^{(28)}^{†} D. Aston,^{(28)}
K.G. Baird,^{(16)} C. Baltay,^{(34)} H.R. Band,^{(33)} M.B. Barakat,^{(34)} G. Baranko,^{(9)}
O. Bardon,^{(15)} T. L. Barklow,^{(28)} G.L. Bashindzhagyan,^{(18)} A.O. Bazarko,^{(10)}
R. Ben-David,^{(34)} A.C. Benvenuti,^{(2)} G.M. Bilei,^{(22)} D. Bisello,^{(21)} G. Blaylock,^{(16)}
J.R. Bogart,^{(28)} B. Bolen,^{(17)} T. Bolton,^{(10)} G.R. Bower,^{(28)} J.E. Brau,^{(20)}
M. Breidenbach,^{(28)} W.M. Bugg,^{(29)} D. Burke,^{(28)} T.H. Burnett,^{(32)}
P.N. Burrows,^{(15)} W. Busza,^{(15)} A. Calcaterra,^{(12)} D.O. Caldwell,^{(5)}
D. Calloway,^{(28)} B. Camanzi,^{(11)} M. Carpinelli,^{(23)} R. Cassell,^{(28)} R. Castaldi,^{(23)}^{(a)}
A. Castro,^{(21)} M. Cavalli-Sforza,^{(6)} A. Chou,^{(28)} E. Church,^{(32)} H.O. Cohn,^{(29)}
J.A. Coller,^{(3)} V. Cook,^{(32)} R. Cotton,^{(4)} R.F. Cowan,^{(15)} D.G. Coyne,^{(6)}
G. Crawford,^{(28)} A. D’Oliveira,^{(7)} C.J.S. Damerell,^{(25)} M. Daoudi,^{(28)}
R. De Sangro,^{(12)} R. Dell’Orso,^{(23)} P.J. Dervan,^{(4)} M. Dima,^{(8)} D.N. Dong,^{(15)}
P.Y.C. Du,^{(29)} R. Dubois,^{(28)} B.I. Eisenstein,^{(13)} R. Elia,^{(28)} E. Etzion,^{(33)}
S. Fahey,^{(9)} D. Falciai,^{(22)} C. Fan,^{(9)} J.P. Fernandez,^{(6)} M.J. Fero,^{(15)} R. Frey,^{(20)}
K. Furuno,^{(20)} T. Gillman,^{(25)} G. Gladding,^{(13)} S. Gonzalez,^{(15)} E.L. Hart,^{(29)}
J.L. Harton,^{(8)} A. Hasan,^{(4)} Y. Hasegawa,^{(30)} K. Hasuko,^{(30)} S. J. Hedges,^{(3)}
S.S. Hertzbach,^{(16)} M.D. Hildreth,^{(28)} J. Huber,^{(20)} M.E. Huffer,^{(28)}
E.W. Hughes,^{(28)} H. Hwang,^{(20)} Y. Iwasaki,^{(30)} D.J. Jackson,^{(25)} P. Jacques,^{(24)}
J. A. Jaros,^{(28)} A.S. Johnson,^{(3)} J.R. Johnson,^{(33)} R.A. Johnson,^{(7)} T. Junk,^{(28)}
R. Kajikawa,^{(19)} M. Kalelkar,^{(24)} H. J. Kang,^{(26)} I. Karliner,^{(13)} H. Kawahara,^{(28)}
H.W. Kendall,^{(15)} Y. D. Kim,^{(26)} M.E. King,^{(28)} R. King,^{(28)} R.R. Kofler,^{(16)}
N.M. Krishna,^{(9)} R.S. Kroeger,^{(17)} J.F. Labs,^{(28)} M. Langston,^{(20)} A. Lath,^{(15)}
J.A. Lauber,^{(9)} D.W.G.S. Leith,^{(28)} V. Lia,^{(15)} M.X. Liu,^{(34)} X. Liu,^{(6)}
M. Loreti,^{(21)} A. Lu,^{(5)} H.L. Lynch,^{(28)} J. Ma,^{(32)} G. Mancinelli,^{(22)} S. Manly,^{(34)}
G. Mantovani,^{(22)} T.W. Markiewicz,^{(28)} T. Maruyama,^{(28)} H. Masuda,^{(28)}
E. Mazzucato,^{(11)} A.K. McKemey,^{(4)} B.T. Meadows,^{(7)} R. Messner,^{(28)}
P.M. Mockett,^{(32)} K.C. Moffeit,^{(28)} T.B. Moore,^{(34)} D. Muller,^{(28)} T. Nagamine,^{(28)}
S. Narita,^{(30)} U. Nauenberg,^{(9)} H. Neal,^{(28)} M. Nussbaum,^{(7)} Y. Ohnishi,^{(19)}
D. Onoprienko,^{(29)} L.S. Osborne,^{(15)} R.S. Panvini,^{(31)} C.H. Park,^{(27)} H. Park,^{(20)}
T.J. Pavel,^{(28)} I. Peruzzi,^{(12)}^{(b)} M. Piccolo,^{(12)} L. Piemontese,^{(11)} E. Pieroni,^{(23)}
K.T. Pitts,^{(20)} R.J. Plano,^{(24)} R. Prepost,^{(33)} C.Y. Prescott,^{(28)} G.D. Punkar,^{(28)}
J. Quigley,^{(15)} B.N. Ratcliff,^{(28)} T.W. Reeves,^{(31)} J. Reidy,^{(17)} P.L. Reinertsen,^{(6)}
P.E. Rensing,^{(28)} L.S. Rochester,^{(28)} P.C. Rowson,^{(10)} J.J. Russell,^{(28)}
O.H. Saxton,^{(28)} T. Schalk,^{(6)} R.H. Schindler,^{(28)} B.A. Schumm,^{(6)}
J. Schwiening,^{(28)} S. Sen,^{(34)} V.V. Serbo,^{(33)} M.H. Shaevitz,^{(10)} J.T. Shank,^{(3)}
G. Shapiro,^{(14)} D.J. Sherden,^{(28)} K.D. Shmakov,^{(29)} C. Simopoulos,^{(28)}
N.B. Sinev,^{(20)} S.R. Smith,^{(28)} M.B. Smy,^{(8)} J.A. Snyder,^{(34)} P. Stamer,^{(24)}
H. Steiner,^{(14)} R. Steiner,^{(1)} M.G. Strauss,^{(16)} D. Su,^{(28)} F. Suekane,^{(30)}
A. Sugiyama,^{(19)} S. Suzuki,^{(19)} M. Swartz,^{(28)} A. Szumilo,^{(32)} T. Takahashi,^{(28)}
F.E. Taylor,^{(15)} E. Torrence,^{(15)} A.I. Trandafir,^{(16)} J.D. Turk,^{(34)} T. Usher,^{(28)}
J. Va’vra,^{(28)} C. Vannini,^{(23)} E. Vella,^{(28)} J.P. Venuti,^{(31)} R. Verdier,^{(15)}
P.G. Verdini,^{(23)} D.L. Wagner,^{(9)} S.R. Wagner,^{(28)} A.P. Waite,^{(28)} S.J. Watts,^{(4)}
A.W. Weidemann,^{(29)} E.R. Weiss,^{(32)} J.S. Whitaker,^{(3)} S.L. White,^{(29)}
F.J. Wickens,^{(25)} D.A. Williams,^{(6)} D.C. Williams,^{(15)} S.H. Williams,^{(28)}
S. Willocq,^{(28)} R.J. Wilson,^{(8)} W.J. Wisniewski,^{(28)} M. Woods,^{(28)} G.B. Word,^{(24)}

J. Wyss,^{(21)} R.K. Yamamoto,^{(15)} J.M. Yamartino,^{(15)} X. Yang,^{(20)} J. Yashima,^{(30)}
S.J. Yellin,^{(5)} C.C. Young,^{(28)} H. Yuta,^{(30)} G. Zapalac,^{(33)} R.W. Zdarko,^{(28)}
and J. Zhou,^{(20)}

(1)Adelphi University, Garden City, New York 11530 (2)INFN Sezione di Bologna, I-40126 Bologna, Italy (3)Boston University, Boston, Massachusetts 02215 (4)Brunel University, Uxbridge, Middlesex UB8 3PH, United Kingdom

(5)University of California at Santa Barbara, Santa Barbara, California 93106

(6)University of California at Santa Cruz, Santa Cruz, California 95064 (7)University of Cincinnati, Cincinnati, Ohio 45221 (8)Colorado State University, Fort Collins, Colorado 80523 (9)University of Colorado, Boulder, Colorado 80309 (10)Columbia University, New York, New York 10027

(11)INFN Sezione di Ferrara and Università di Ferrara, I-44100 Ferrara, Italy

(12)INFN Lab. Nazionali di Frascati, I-00044 Frascati, Italy (13)University of Illinois, Urbana, Illinois 61801

(14)E.O. Lawrence Berkeley Laboratory, University of California, Berkeley,

California 94720

(15)Massachusetts Institute of Technology, Cambridge, Massachusetts 02139

(16)University of Massachusetts, Amherst, Massachusetts 01003 (17)University of Mississippi, University, Mississippi 38677

(18)Moscow State University, Institute of Nuclear Physics 119899 Moscow, Russia

(19)Nagoya University, Chikusa-ku, Nagoya 464 Japan (20)University of Oregon, Eugene, Oregon 97403

(21)INFN Sezione di Padova and Università di Padova, I-35100 Padova, Italy (22)INFN Sezione di Perugia and Università di Perugia, I-06100 Perugia, Italy

(23)INFN Sezione di Pisa and Università di Pisa, I-56100 Pisa, Italy (24)Rutgers University, Piscataway, New Jersey 08855

(25)Rutherford Appleton Laboratory, Chilton, Didcot, Oxon OX11 0QX United

Kingdom (26)Sogang University, Seoul, Korea (27)Soongsil University, Seoul, Korea 156-743

(28)Stanford Linear Accelerator Center, Stanford University, Stanford, California

94309 (29)University of Tennessee, Knoxville, Tennessee 37996 (30)Tohoku University, Sendai 980 Japan (31)Vanderbilt University, Nashville, Tennessee 37235 (32)University of Washington, Seattle, Washington 98195 (33)University of Wisconsin, Madison, Wisconsin 53706 (34)Yale University, New Haven, Connecticut 06511 †Deceased (a)Also at the Università di Genova (b)Also at the Università di Perugia

This work was supported by Department of Energy contracts: DE-FG02-91ER40676 (BU), DE-FG03-91ER40618 (UCSB), DE-FG03-92ER40689 (UCSC), DE-FG03-93ER40788 (CSU), DE-FG02-91ER40672 (Colorado), DE-FG02-91ER40677 (Illinois), DE-AC03-76SF00098 (LBL), DE-FG02-92ER40715 (Massachusetts), DE-FC02-94ER40818 (MIT), DE-FG03-96ER40969 (Oregon), DE-AC03-76SF00515 (SLAC), DE-FG05-91ER40627 (Tennessee), DE-FG02-95ER40896 (Wisconsin), DE-FG02-92ER40704 (Yale); National Science Foundation grants: PHY-91-13428 (UCSC), PHY-89-21320 (Columbia), PHY-92-04239 (Cincinnati), PHY-95-10439 (Rutgers), PHY-88-19316 (Vanderbilt), PHY-92-03212 (Washington); The UK Particle Physics and Astronomy Research Council (Brunel and RAL); The Istituto Nazionale di Fisica Nucleare of Italy (Bologna, Ferrara, Frascati, Pisa, Padova, Perugia); The Japan-US Cooperative Research Project on High Energy Physics (Nagoya, Tohoku); The Korea Science and Engineering Foundation (Soongsil).